OpenStudy (anonymous):

At what value would 5x+12 and 2x+14 be congruent

OpenStudy (imstuck):

If those are congruent, that means that they are equal to one another. So set them equal to one another and solve for x.

OpenStudy (imstuck):

5x+12=2x+14 3x=2 x=2/3 For the value of x = 2/3, the expressions are congruent.

OpenStudy (anonymous):

Thank you